# arXiv:2602.23189v2[math.AP] 02 Mar 2026

### v2

@@ 3.1 A key $L^{2}$-estimate for the partial densities
Assume that there exists a constant $c>0$ such that $R_{+}^{\varepsilon}>c$. Therefore, we have
$\lvert\alpha_{+}^{\varepsilon}-\alpha_{+}\rvert\leqslant C\lvert R_{+}^{\varepsilon}-R_{+}\rvert+C\lvert R_{-}^{\varepsilon}-R_{-}\rvert$. (3.3)
@@
$\lvert a_{1}-a_{2}\rvert\leqslant C\lvert d_{1}-d_{2}\rvert$. (3.6)
@@
We then define the function $g$ by setting $a_{i}=g(d_{i})$ for all $i$. The monotonicity of $F_{d_{i}}$ with respect to $x$ implies that $g$ is decreasing with respect to $d_{i}$.
First, consider the case where $d_{1}<\frac{\underline{c}}{2}$ or $d_{1}>2\overline{c}$. Since $d_{2}\in[\underline{c},\overline{c}]$, there exists a constant $C>0$ sufficiently large such that
@@
Now consider the case where $d_{1}\in[\frac{\underline{c}}{2},2\overline{c}]$. Since the function $g$ is decreasing, it follows that $a_{1}\in[g(2\overline{c}),g(\frac{\underline{c}}{2})]$. We then introduce the compact set $K=[\frac{\underline{c}}{2},2\overline{c}]\times[g(2\overline{c}),g(\frac{\underline{c}}{2})]$ and and define the function $F:K\to\mathbb{R}$ by $F(d,x)=dx^{\gamma}+x-1$. The function $F$ is of class $C^{\infty}$ on $K$.
For any $(d_{i},x_{i})\in K$ satisfying $F(d_{i},x_{i})=0$ and $\partial_{x}F(d_{i},x_{i})\neq 0$, the Implicit Function Theorem ensures the existence of an open neighbourhood $U_{i,\mathrm{loc}}$ of $d_{i}$ and a unique function $g_{i,\mathrm{loc}}:U_{i,\mathrm{loc}}\to\mathbb{R}$ such that
$x_{i}=g_{i,\mathrm{loc}}(d_{i})\quad\text{and}\quad F\bigl(d,g_{i,\mathrm{loc}}(d)\bigr)=0\quad\text{for all }d\in U_{i,\mathrm{loc}}.$
Since the solution $(d_{i},x_{i})$ of the equation $F(d_{i},x)=0$ is unique, it follows from the uniqueness that
$g_{i,\mathrm{loc}}=g.$
@@
$g_{i,\mathrm{loc}}^{\prime}(d)=-\,\frac{\partial_{d}F(d,g(d))}{\partial_{x}F(d,g(d))}=-\,\frac{x^{\gamma}}{\gamma dx^{\gamma-1}+1},\qquad\text{with }x=g(d).$
@@
$\left|g_{i,\mathrm{loc}}^{\prime}(d)\right|\leqslant L\quad\text{for all }d\in\Bigl[\frac{\underline{c}}{2},2\overline{c}\Bigr].$
@@
Having established Lemma 3.2, we can now complete the proof of the main result of this section.
@@
If we denote by $d_{1}=\frac{R_{-}^{\varepsilon}}{(R_{+}^{\varepsilon})^{\gamma}}$ and $d_{2}=\frac{R_{-}}{R_{+}^{\gamma}}$, the hypothesis $d_{2}\in[\underline{c},\overline{c}]$ is verified because $\alpha_{\pm}\in[\underline{a},\overline{a}]\subset(0,1)$, see Remark 1.4, and we can apply Lemma 3.2. It follows that we obtain a unique solution $(\alpha_{+}^{\varepsilon},\alpha_{+})\in(0,1)$ such that equations (3.8) and (3.9) are satisfied and that there exists a constant $C$ such that
@@
By adding and substracting $\frac{R_{-}}{(R_{+}^{\varepsilon})^{\gamma}}$ and using the assumption $R_{+}^{\varepsilon}>c$, we obtain
@@
Finally, using the mean value theorem for the function $f(x)=\frac{1}{x^{\gamma}}$ defined for $x\in\mathbb{R}_{+},x>\underline{x}$, we can conclude the proof.
As a consequence of the previous lemma, we manage to estimate the $L^{2}$ norm of $R_{\pm}^{\varepsilon}-R_{\pm}$ by the entropy and the $L^{1}$ norm of $R_{\pm}^{\varepsilon}-R_{\pm}$, such a result is given in the following proposition.
Proposition 3.3.
Under the assumptions of Theorem 1.6. Let $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ a weak solution of System (LABEL:system_eps), in the sense of Definition 2.1, and $(\rho_{+},\rho_{-},\alpha_{+},u)$ a strong solution of System (1.5), as in Theorem 1.3, there exists $C>0$ depending only on $\underline{a},\overline{a},C_{0}$ such that
$\|R_{\pm}^{\varepsilon}-R_{\pm}\|_{L^{2}}^{2}\leqslant C\varepsilon^{2}{\mathcal{E}}_{1}+C\|R_{+}^{\varepsilon}-R_{+}\|_{L^{1}}+C\|R_{-}^{\varepsilon}-R_{-}\|_{L^{1}}.$
Proof.
We have
$\int_{\Omega}\lvert R_{\pm}^{\varepsilon}-R_{\pm}\rvert^{2}$ $=\int_{\Omega}\lvert\alpha_{\pm}^{\varepsilon}(\rho_{\pm}^{\varepsilon}-\rho_{\pm})+(\alpha_{\pm}^{\varepsilon}-\alpha_{\pm})\rho_{\pm}\rvert^{2}$
$\leqslant C\int_{\Omega}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}+C\int_{\Omega}\rho_{\pm}^{2}\lvert\alpha_{\pm}^{\varepsilon}-\alpha_{\pm}\rvert^{2}.$
The first term is controlled by the relative entropy ${\mathcal{E}}_{1}$ through Inequalities (3.1)–(3.2):
$\int_{\Omega}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}$ $=\int_{\{\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert\leqslant 1\}}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}+\int_{\{\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert>1\}}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{\gamma_{\pm}}$
$\leqslant C\varepsilon^{2}{\mathcal{E}}_{1}.$ (3.10)
where we have used $\gamma_{+},\gamma_{-}\geqslant 2$.
For the second term, we want to show that
$\int_{\Omega}\rho_{\pm}^{2}\lvert\alpha_{\pm}^{\varepsilon}-\alpha_{\pm}\rvert^{2}\leqslant C\varepsilon^{2}{\mathcal{E}}_{1}+C\sum_{i=+,-}\int_{\Omega}\lvert R_{i}^{\varepsilon}-R_{i}\rvert$. (3.11)
@@
    which, by using (3.10), can be estimated by the right-hand side terms of (3.11).
@@
This ends the proof of (3.11) and of this proposition. ∎
3.2 Csiszár-Kullback-Pinsker-type inequality
We will now present an inequality close to the Csiszar-Kullback-Pinsker inequality.
Proposition 3.4.
Let $f$ a nonnegative and $g$ a positive real functions in $L^{1}$ such that $\int f=\int g$. Then
$\|f-g\|_{L^{1}}\leqslant\int f\ln_{+}\frac{f}{g}$

### v3

@@ 5 Strong convergences
Assume that there exists a constant $c>0$ such that $R_{+}^{\varepsilon}>c$. Therefore, there exists $C>0$ such that
$\lvert\alpha_{+}^{\varepsilon}-\alpha_{+}\rvert\leqslant C\lvert R_{+}^{\varepsilon}-R_{+}\rvert+C\lvert R_{-}^{\varepsilon}-R_{-}\rvert$.
@@
$\lvert a_{1}-a_{2}\rvert\leqslant C\lvert d_{1}-d_{2}\rvert$.
@@
We then define the function $g:[\underline{c},\overline{c}]\to(0,1)$ by setting $a=g(d)$ such that $F_{d}(a)=0$. The monotonicity of $F_{d}$ with respect to $x$ implies that $g$ is decreasing with respect to $d$.
First, consider the case where $d_{1}<\underline{c}/2$ or $d_{1}>2\overline{c}$. Since $d_{2}\in[\underline{c},\overline{c}]$, there exists a constant $C>0$ sufficiently large such that
@@
Now consider the case where $d_{1}\in[\underline{c}/2,2\overline{c}]$. Since the function $g$ is decreasing, it follows that $a_{1}\in[g(2\overline{c}),g(\frac{\underline{c}}{2})]$. We then introduce the compact set $K=[\frac{\underline{c}}{2},2\overline{c}]\times[g(2\overline{c}),g(\frac{\underline{c}}{2})]$ and and define the function $F:K\to\mathbb{R}$ by $F(d,x)=dx^{\gamma}+x-1$. The function $F$ is of class $C^{\infty}$ on $K$.
For any $(d_{0},x_{0})\in K$ satisfying $F(d_{0},x_{0})=0$, as $\partial_{x}F(d_{0},x_{0})\neq 0$, the Implicit Function Theorem ensures the existence of an open neighbourhood $U_{0,\mathrm{loc}}$ of $d_{0}$ and a unique function $g_{0,\mathrm{loc}}:U_{0,\mathrm{loc}}\to\mathbb{R}$ such that
$x=g_{0,\mathrm{loc}}(d)\quad\text{and}\quad F\bigl(d,g_{0,\mathrm{loc}}(d)\bigr)=0\quad\text{for all }d\in U_{0,\mathrm{loc}}.$
Since the solution $(d,x)$ of the equation $F(d,x)=0$ is unique, it follows that
$g_{0,\mathrm{loc}}=g.$
@@
$g_{0,\mathrm{loc}}^{\prime}(d)=-\,\frac{\partial_{d}F(d,g(d))}{\partial_{x}F(d,g(d))}=-\,\frac{x^{\gamma}}{\gamma dx^{\gamma-1}+1},\qquad\text{with }x=g(d).$
@@
$\left|g^{\prime}(d)\right|\leqslant L\quad\text{for all }d\in\Bigl[\frac{\underline{c}}{2},2\overline{c}\Bigr].$
@@
Having established Lemma 5.2, we can now complete the proof of Lemma 5.1.
@@
If we denote by $d_{1}=\frac{R_{-}^{\varepsilon}}{(R_{+}^{\varepsilon})^{\gamma}}$ and $d_{2}=\frac{R_{-}}{R_{+}^{\gamma}}$, the hypothesis $d_{2}\in[\underline{c},\overline{c}]$ is verified because $\alpha_{\pm}\in[\underline{a},\overline{a}]\subset(0,1)$, see Remark 2.4, and we can apply Lemma 5.2. It follows that we obtain a unique solution $(\alpha_{+}^{\varepsilon},\alpha_{+})\in(0,1)$ such that Equations (5.1)-(5.2) are satisfied and that there exists a constant $C$ such that
@@
By adding and subtracting $\frac{R_{-}}{(R_{+}^{\varepsilon})^{\gamma}}$ and using the assumption $R_{+}^{\varepsilon}>c$, we obtain
@@
Finally, using the mean value theorem for the function $f(x)=\frac{1}{x^{\gamma}}$ defined for $x\in[\min(c,R_{+,0}),+\infty)$, we can conclude the proof. ∎
As a consequence of Lemma 5.1, we want to show that
$\int_{\Omega}\rho_{\pm}^{2}\lvert\alpha_{\pm}^{\varepsilon}-\alpha_{\pm}\rvert^{2}\leqslant C\varepsilon^{2}{\mathcal{E}}^{\varepsilon}_{1}+C\sum_{i=+,-}\int_{\Omega}\lvert R_{i}^{\varepsilon}-R_{i}\rvert$. (5.3)
@@
    which is controlled by the relative entropy ${\mathcal{E}}^{\varepsilon}_{1}$ (1.7) through Inequality (3.1):
    $\int_{\Omega}\alpha_{\pm}^{\varepsilon}\lvert\rho_{\pm}^{\varepsilon}-\rho_{\pm}\rvert^{2}\leqslant C\int_{\Omega}\alpha_{\pm}^{\varepsilon}H_{\pm}(\rho_{\pm}^{\varepsilon}|\rho_{\pm})\leqslant C\varepsilon^{2}{\mathcal{E}}^{\varepsilon}_{1}.$
    where we have used $\gamma_{+},\gamma_{-}\geqslant 2$.
@@
where we also have used $\lvert\alpha_{\pm}^{\varepsilon}-\alpha_{\pm}\rvert\leqslant 2$. This ends the proof of (5.3).
By the uniform estimates of ${\mathcal{E}}_{1}^{\varepsilon}(t)$, we finish the proof of Theorem 1.3 by noticing that (5.3) reads
$\|(\alpha_{i}^{\varepsilon}-\alpha_{i})(t,\cdot)\|_{L^{2}(\Omega)}^{2}\leqslant C\varepsilon^{2}+C\sum_{i=+,-}\|(R_{i}^{\varepsilon}-R_{i})(t,\cdot)\|_{L^{2}(\Omega)}\to 0\quad\text{in }L^{q}(0,T)$
for any $q\in[1,\infty)$.
Acknowledgements. The author would like to thank Didier Bresch and Christophe Lacave for their many tips on writing this text. She also gratefully acknowledges the University Savoie Mont Blanc and the Agence Nationale pour la Recherche (ANR) for her PhD position fellowship linked to the CPJ Anamod managed by Christophe Lacave. She also wants to thank for some support of the ANR under France 2030 bearing the reference ANR-23-EXMA-004 (Complexflows project) related to the PEPR Maths-ViVEs. The author wants also to gratefully acknowledge the partial support by the ANR grant ANR-23-CE40-0014-01 (ANR Bourgeons).
Appendix A Entropy computations
In this appendix, we derive entropy relations for general velocity fields $u$, extending the computations presented in the main text (see Section 3.1) and providing a framework that could be applied to stability analyses or weak-strong uniqueness studies. For that, we consider two solutions of System (LABEL:system) (without nondimensionalization and the Mach number): a weak solution still denoted by indices $\varepsilon$, $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$, and a sufficiently regular solution $(\rho_{+},\rho_{-},\alpha_{+},u)$. We define the following quantities:
  • the energy
    $E(t):=\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u^{\varepsilon}|^{2}\Big)(t,x)\,\mathrm{d}x+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}(\rho_{i}^{\varepsilon})^{\gamma_{i}}(t,x)}{\gamma_{i}-1}\,\mathrm{d}x,$
  • the modulated entropy
    ${\mathcal{E}}_{1}(t)$ $={\mathcal{E}}_{1}\Big(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon}|\rho_{+},\rho_{-},\alpha_{+},u\Big)$
    $=\frac{1}{2}\int_{\Omega}\rho^{\varepsilon}(t,x)|u^{\varepsilon}-u|^{2}(t,x)\,\mathrm{d}x+\sum_{i=+,-}\int_{\Omega}\alpha_{i}^{\varepsilon}(t,x)H(\rho_{i}^{\varepsilon}|\rho_{i})(t,x)\,\mathrm{d}x.$
Proposition A.1.
If $(\rho^{\varepsilon}_{+},\rho^{\varepsilon}_{-},\alpha^{\varepsilon}_{+},u^{\varepsilon})$ is a weak solution of System (LABEL:system), defined as in Definition 2.1, then for all $(\rho_{+},\rho_{-},\alpha_{+},u)$ sufficiently regular solutions of the same System (LABEL:system), we have
${\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)\leqslant$ $-\mu\int_{0}^{t}\!\!\!\int_{\Omega}\left(\nabla(u^{\varepsilon}-u)\right)^{2}-(\lambda+\mu)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}(u^{\varepsilon}-u)\right)^{2}\,\mathrm{d}x\,\mathrm{d}s$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(\partial_{t}u+u\cdot\nabla u)\cdot(u-u^{\varepsilon})\,\mathrm{d}x\,\mathrm{d}s$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\left(\alpha_{+}^{\varepsilon}\rho_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon}\rho_{-}^{\varepsilon}\right)(u^{\varepsilon}-u)\cdot\nabla u\cdot(u-u^{\varepsilon})\,\mathrm{d}x\,\mathrm{d}s$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\Big(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-})-\left(\alpha_{+}^{\varepsilon}p_{+}\left(\rho_{+}^{\varepsilon}\right)+\alpha_{-}^{\varepsilon}p_{-}\left(\rho_{-}^{\varepsilon}\right)\right)\Big)\operatorname{div}u\,\mathrm{d}x\,\mathrm{d}s$
$+\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}\,\mathrm{d}x\,\mathrm{d}s$
$+\sum_{i=+,-}\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}u-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}\,\mathrm{d}x\,\mathrm{d}s$
$-\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u:\nabla(u^{\varepsilon}-u)-(\lambda+\mu)\,\mathrm{d}x\,\mathrm{d}s\int_{0}^{t}\!\!\!\int_{\Omega}\operatorname{div}u\operatorname{div}(u^{\varepsilon}-u)\,\mathrm{d}x\,\mathrm{d}s.$
Remark A.2.
The equation satisfied by the strong solution $(\rho_{\pm},\alpha_{+},u)$ is not written explicitly in compact form in the right-hand side above. Instead, its contribution appears through the various decomposed terms, which correspond to the expansion of the momentum equation.
Proof.
We have
${\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)=$ $E(t)-E(0)$ (A.1)
$+\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(t,\cdot)-\frac{1}{2}\int_{\Omega}\Big(\rho^{\varepsilon}|u|^{2}\Big)(0,\cdot)$
$-\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(t,\cdot)+\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(0,\cdot)$
$-\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(t,\cdot)$
$+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(0,\cdot).$
Let us now use the momentum equation by taking $\phi=u$, we get:
$\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(t,\cdot)+\int_{\Omega}\Big(\rho^{\varepsilon}u^{\varepsilon}\cdot u\Big)(0,\cdot)=-\int_{0}^{t}\!\!\!\int_{\Omega}\rho^{\varepsilon}u^{\varepsilon}\cdot(\partial_{t}u+(u^{\varepsilon}\cdot\nabla)u)\\ -\int_{0}^{t}\!\!\!\int_{\Omega}\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u^{\varepsilon}:\nabla u+(\mu+\lambda)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}u^{\varepsilon}\right)(\operatorname{div}u),$
where we can replace
$-\int_{0}^{t}\!\!\!\int_{\Omega}\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u$ $=-\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon})\rho_{+}^{\varepsilon\gamma_{+}}\operatorname{div}u$
$=-\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+}^{\varepsilon})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-}^{\varepsilon}))\operatorname{div}u.$
We can also rewrite
$\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i}\Big)=\sum_{i=+,-}\int_{\Omega}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}}=\int_{\Omega}\rho_{+}^{\gamma_{+}}$
hence
$\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i}\Big)(t,\cdot)+\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}-\gamma_{i}\rho_{i}^{\gamma_{i}-1}\rho_{i})\Big)(0,\cdot)\\ =\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}+\alpha_{-}^{\varepsilon})\partial_{t}\rho_{+}^{\gamma_{+}}=\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\alpha_{i}^{\varepsilon}\gamma_{i}\rho_{i}^{\gamma-1}\partial_{t}\rho_{i}$
$-\int_{\Omega}\frac{\gamma_{i}\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\rho_{i}^{\gamma_{i}-1}\rho_{i}^{\varepsilon}(t,\cdot)+\int_{\Omega}\frac{\gamma_{i}\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\rho_{i}^{\gamma_{i}-1}\rho_{i}^{\varepsilon}(0,\cdot)$ $=-\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}\Big(\partial_{t}\rho_{i}+(u^{\varepsilon}\cdot\nabla)\rho_{i}\Big).$
We then have obtained
$-\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(t,\cdot)$
$+\sum_{i=+,-}\int_{\Omega}\frac{\alpha_{i}^{\varepsilon}}{(\gamma_{i}-1)}\Big(\rho_{i}^{\gamma_{i}}+\gamma_{i}\rho_{i}^{\gamma_{i}-1}(\rho_{i}^{\varepsilon}-\rho_{i})\Big)(0,\cdot)$ $=\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}$
$+\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}u-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}$
$-\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}\alpha_{i}\rho_{i}u\cdot\nabla\rho_{i}$
$-\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}\alpha_{i}\rho_{i}u\cdot\nabla\rho_{i}=$ $-\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\alpha_{i}u\cdot\nabla p_{i}(\rho_{i})=-\int_{0}^{t}\!\!\!\int_{\Omega}u\cdot\nabla p_{+}(\rho_{+})$
$=$ $\int_{0}^{t}\!\!\!\int_{\Omega}p_{+}(\rho_{+})\operatorname{div}u$
$=$ $\int_{0}^{t}\!\!\!\int_{\Omega}(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-}))\operatorname{div}u$
where we have used that $u$ satisfies the Dirichlet boundary condition.
Therefore, using the energy estimate (2.2) satisfied by $(\rho_{+}^{\varepsilon},\rho_{-}^{\varepsilon},\alpha_{+}^{\varepsilon},u^{\varepsilon})$ and plugging all the previous computations into (A.1), we obtain:
${\mathcal{E}}_{1}(t)-{\mathcal{E}}_{1}(0)$ $\leqslant-\mu\int_{0}^{t}\!\!\!\int_{\Omega}\left(\nabla u^{\varepsilon}\right)^{2}\,\mathrm{d}x\,\mathrm{d}t-\left(\lambda+\mu\right)\int_{0}^{t}\!\!\!\int_{\Omega}(\operatorname{div}u^{\varepsilon})^{2}\,\mathrm{d}x\,\mathrm{d}t$
$+\int_{0}^{t}\!\!\!\int_{\Omega}\Big(\alpha_{+}^{\varepsilon}p_{+}(\rho_{+})+\alpha_{-}^{\varepsilon}p_{-}(\rho_{-})-\left(\alpha_{+}^{\varepsilon}p_{+}\left(\rho_{+}^{\varepsilon}\right)+\alpha_{-}^{\varepsilon}p_{-}\left(\rho_{-}^{\varepsilon}\right)\right)\Big)\operatorname{div}u$
$+\sum_{i=+,-}\int_{0}^{t}\!\!\!\int_{\Omega}\gamma_{i}\alpha_{i}^{\varepsilon}\rho_{i}^{\gamma_{i}-2}(\rho_{i}-\rho_{i}^{\varepsilon})\partial_{t}\rho_{i}+\sum_{i=+,-}\int_{\Omega}\gamma_{i}\rho_{i}^{\gamma_{i}-2}(\alpha_{i}\rho_{i}-\alpha_{i}^{\varepsilon}\rho_{i}^{\varepsilon}u^{\varepsilon})\cdot\nabla\rho_{i}$
$+\mu\int_{0}^{t}\!\!\!\int_{\Omega}\nabla u^{\varepsilon}:\nabla u+(\mu+\lambda)\int_{0}^{t}\!\!\!\int_{\Omega}\left(\operatorname{div}u^{\varepsilon}\right)(\operatorname{div}u).$
$\nabla u^{\varepsilon}:\nabla(u^{\varepsilon}-u)=-\left(\nabla(u^{\varepsilon}-u)\right)^{2}-\nabla u:\nabla(u^{\varepsilon}-u),$